OpenStudy (anonymous):

Find the equivalent exponential expression. [(-4)^5]^3 A.(-4)^-2 B.(-4)^8 C.(-4)^2 D.(-4)^15

OpenStudy (anonymous):

This is a problem that deals with the properties of exponents. When a value that is raised to a certain power is raised to another power, they multiply. Basically, \[(x^{a})^b=x^{ab}\]so in this problem, simply multiply the 5 and 3 to give you \[(-4)^{15}\]

OpenStudy (imstuck):

Just like before, this is equivalent to \[(-4^{5})^{3}=(-4)^{15}=-1073741824\]

OpenStudy (anonymous):

so what about this 12^21?

OpenStudy (anonymous):

That's just a huge number.\[12^{21}=4.6\times 10^{22}\]

OpenStudy (anonymous):

ah here are the available options for answers : A. (12)^20 B.(12^8)^-12 C.(12^7)^3 D.(12^7)^14

OpenStudy (anonymous):

Oh I see. So it's the same concept as before. Look for the 2 exponents that multiply to be 21. Looking through all the answers, C is the most logical because\[(12^{7})^3=12^{21}\]

OpenStudy (anonymous):

Wow okay thanks very much I pretty much got all the rest after that one, it really is just that easy. again thank you!

OpenStudy (anonymous):

No problem! Just make sure to know all the different properties of exponents and you'll be able to solve them a lot easier!